3r-5s=21
5r+3s=33
solve by elimination method

Answer by stanbon(75887)

3r-5s=21
5r+3s=33
solve by elimination method
-------------------
Multiply 1st by 5
Multiply 2nd by 3
-----------
15r - 25s = 105
15r + 9s = 99
-----
Subtract 1st from 2nd to get:
34s = -6
s = -3/17
==================
Substitute into 3r-5s = 21 to solve for r:
3r -5(-3/17) = 21
3r = 21 -15/17
r = 7 - 5/17
r = 114/17
===================
Cheers,
Stan H.